OpenStudy (anonymous):

Solve R = ts^2 + 3 for s

OpenStudy (anonymous):

R-3 = ts^2 ts^2 = R-3 s^2 = (R-3)/t s = root under((R-3)/t)

OpenStudy (anonymous):

thanx but I'm kind of confused where the root goes?

OpenStudy (anonymous):

its whole root under that means...whole expression goes under root sign

OpenStudy (anonymous):

Does it go over the whole side of the equation except for s?

OpenStudy (anonymous):

oh ok thanx I get it now:)

OpenStudy (anonymous):

\[\sqrt{(R-3)/t}\]